A carbon resistor is to be used as a thermometer. On a winter day when the temperature is 4.00 ∘C, the resistance of the carbon resistor is 217.7 Ω . What is the temperature on a spring day when the resistance is 215.4 Ω ? Take the temperature coefficient of resistivity for carbon to be α = −5.00×10−4 C−1 .?
Answer #1

Resistance of resistor at 4°C=R1=217.7 Ohms

New resistance=R2=215.4 Ohms

Let the Temperature at which the resistance will be 215.4 Ohms be T

Coefficient of resistivity = -5.00*10^-4 /°C

R2=R1*(1+(-5*10^-4*(T-217.7))

215.4=217.7*(1+(-5*10^-4)*(T-217.7))

T=196.57 °C

Therefore Temperature at which resistance will be 215.5 ohms is 196.57°C